OpenStudy (anonymous):

Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). 2y/2 -4x/2 +5/2 m=2 y=-4x+20 ?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

4x + 2y = 5 i think your slope is wrong... subtract 4x on both sides: 2y = -4x + 5 divide everything by 2: y = -2x+5 so your slope is m=-2, for a parallel line. Now use point-slope form again, y-y1 = m(x-x1) with the point (-3,5)

OpenStudy (anonymous):

y-5=-2x+6 then -5 on both sides y=-2x+1

OpenStudy (agent0smith):

very good! but i think you missed a negative... y-5 = -2(x-(-3))

OpenStudy (anonymous):

oh yeah lol, y=-2x-1

OpenStudy (agent0smith):

:D

OpenStudy (anonymous):

I have on now with perpendicular, would I do the same steps?

OpenStudy (agent0smith):

basically BUT remember you have to get the negative reciprocal of the original line's slope.